OpenStudy (anonymous):

Please help me out. :) An unknown charge and a charge of 60 nC are 2m apart. The electric field intensity is O at a point 0.8m from the unknown charge on the line connecting the two charges. Find the unknown charge.

OpenStudy (anonymous):

The electric field is defined as: \(\vec{E} = \vec{F}/\left| q \right|\)

OpenStudy (anonymous):

So, if the electric field is zero, then the force on that point is zero

OpenStudy (anonymous):

\(\sum \vec{E} = \vec{0}\) \[\frac{k*q_1}{(2+0.8)^2} + \frac{k*q_2}{0.8^2} = 0\], solve for \(q_2\), where \(q_1=60nC\)

OpenStudy (anonymous):

and \(k = 8.99\times10^9 N.m^2/C^2\)

OpenStudy (anonymous):

I'll just give my final answer... Unknown Charge = 2.667 x 10^-8 N/C Is it right?